Q15. Consulting income at Kate Walsh Associates for the period January – May has been as follows:

Month

Income

January

7

February

8

March

9

April

6

May

5

Use 3-month weighted moving average to forecast June’s income with weights of 3 for the most recent period, 2 for the second most recent period, and 1 for the third most recent period.
Answer #1

3-month weighted moving average to forecast June’s income =(w1*x1+w2*x2+w3*x3)/(w1+w2+w3)

=(3*5+2*6+1*9)/(3+2+1) =6